OpenStudy (anonymous):

evaluate √x+7 when x = 9. write the answer in simplified radical form 10 16 4 8

OpenStudy (anonymous):

@mathmale

OpenStudy (abdullahm):

Hey there!

OpenStudy (anonymous):

hello

OpenStudy (abdullahm):

Is it the square root of x or the square root of (x + 7) ?

OpenStudy (anonymous):

the square root of (x + 7)

OpenStudy (anonymous):

i think the answer is 4

OpenStudy (abdullahm):

Ok, great! So, first of all, we need to plug in x = 9 What is 9 + 7 = ?

OpenStudy (anonymous):

16

OpenStudy (abdullahm):

And so we have: sqrt(9 + 7) = sqrt (16) = ? What is the square root of 16? Hint: What times itself will give you 16?

OpenStudy (anonymous):

4 x 4

OpenStudy (abdullahm):

So sqrt (16) = 4 Are all of the answer choices under square roots or no?

OpenStudy (anonymous):

wait sorry i didnt get your question?

OpenStudy (abdullahm):

Are all of those numbers in your answer choices under a \(\sqrt{}\)? Or are the answers choices just a number like 4?

OpenStudy (anonymous):

there just numbers like 4

OpenStudy (abdullahm):

Then your answer is just 4. :)

OpenStudy (anonymous):

thanks! can you help me with 1 more question

OpenStudy (abdullahm):

Sure!

OpenStudy (anonymous):

evaluate √24 - x when x = 6. write the answer in simplified radical form. √2 3 3√2 9

OpenStudy (anonymous):

i think it's 9 because 9 cant be simplified

OpenStudy (abdullahm):

So is it sqrt (24 - x) ?

OpenStudy (anonymous):

yep

OpenStudy (abdullahm):

What is 24 - 6 = ?

OpenStudy (anonymous):

18

OpenStudy (abdullahm):

Now we need to simplify sqrt 18. sqrt 18 = sqrt (9 * 2) = sqrt(9) * sqrt (2) What is sqrt 9 = ?

OpenStudy (anonymous):

2?

OpenStudy (abdullahm):

Does 2 * 2 = 9?

OpenStudy (anonymous):

no 3 * 3 does so the sqrt of 9 is 3

OpenStudy (abdullahm):

Yes, so now we have sqrt(9) * sqrt (2) = 3 * sqrt (2) = 3sqrt(2)

OpenStudy (abdullahm):

Do you understand how I simplified square root 18?

OpenStudy (anonymous):

yes i do thanks

OpenStudy (anonymous):

one more question? please :(

OpenStudy (abdullahm):

It was my pleasure. :)

OpenStudy (abdullahm):

Sure!

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

evaluate √16 - x when x = 8. write the answer as a decimal 4 1.41 2 2.83

OpenStudy (anonymous):

round to the nearest hundreds if necessary**

OpenStudy (anonymous):

i think its D

OpenStudy (abdullahm):

sqrt (16 - x), right? The parenthesis make a big difference.

OpenStudy (anonymous):

yes

OpenStudy (abdullahm):

It is D. You are correct. :)

OpenStudy (anonymous):

yay thanks for your help!

OpenStudy (abdullahm):

I'm glad that I was able to be of some help. :)